# Random local access for sampling $k$-SAT solutions

### v1

@@ 3 A random local access algorithm for $k$-CNF solutions
Input: CNF formula $\Phi=(V,\mathcal{C})$, $v\in V$, partial assignment $\sigma$
      Output: random assignment $\sigma(v)\in\{0,1\}$
1:Initialize $\mathcal{S}=\{C^{\sigma}\in\mathcal{C}^{\sigma}\mid v\in\text{vbl}(C^{\sigma})\}$ as a stack
@@
6:              $\sigma(w)\leftarrow\mathsf{MarginSample}(\Phi,\sigma,w)$
@@
12:              Push $\{(C^{\prime})^{\sigma}\mid C^{\prime}\in\mathcal{C},w\in\operatorname{vbl}((C^{\prime})^{\sigma}),(C^{\prime})^{\sigma}\neq 1\}$ onto $\mathcal{S}$               return $\Phi_{v}^{\sigma}$
Algorithm 1 contains several subalgorithms that are formally given above. We describe more colloquially how these algorithms work and are interconnected below. To sample the assignment of some $v\in V$, we first use $\mathsf{IsMarked}(\cdot)$ as in Theorem 2.12 to determine whether $v$ lies in a given $\alpha$-marking $\mathcal{M}$. The next steps of our sampling depend on this boolean outcome.
At the beginning of Algorithm 1, we have a target variable $v$ we wish to sample and initialize partial assignment $\sigma=\emptyset$. We first compute $\mathsf{IsMarked}(v)$. If $\mathsf{IsMarked}(v)=1$, then we perform $\mathsf{MarginSample}(\Phi,\sigma,v)$ (Algorithm 2) to sample the assignment of $v$. With probability $2\theta$ we uniformly at random select $v$ from $\{0,1\}$, and with probability $1-2\theta$ we sample from the overflow distribution $\mathcal{D}:=\frac{\mu_{v}^{\sigma}-\theta}{1-2\theta}$ by calling $\mathsf{MarginOverflow}(\Phi,\sigma_{v\leftarrow\star},v)$ (Algorithm 3). The notation $v\leftarrow\star$ means that we will temporarily mark $v$ as an accessed by unassigned variable (2.2) and sample its value in the end.
Going into $\mathsf{MarginOverflow}(\Phi,\sigma_{v\leftarrow\star},v)$, we will use $\mathsf{NextVar}(\cdot)$ to determine the smallest unaccessed marked neighbor of $v$ (call it $w$). Again, with probability $2\theta$ we randomly select $w$ from $\{0,1\}$, and with probability $1-2\theta$, we sample $w$ from the overflow distribution $\mathsf{MarginOverflow}(\Phi,\sigma_{w\leftarrow\star},u)$. We will show in Proposition 4.4 that, with high probability, this process terminates after a polylogarithmic number of $\mathsf{MarginOverflow}$ iterations. After that we compute $\Phi^{\sigma}_{v}$, the connected component that contains $v$ under the current partial assignment $\sigma$. Finally we sample from $\Phi^{\sigma}_{v}$ using a uniform sampling algorithm (Theorem 4.3) and output its $v$ coordiante, concluding the procedure $\mathsf{MarginSample}(\Phi,\emptyset,v)$.
Otherwise, if $v$ is not marked in $\mathcal{M}$, then we compute $\mathsf{Conn}(\Phi,\emptyset,v)$ (Algorithm 4), the connected component in $\Phi^{\sigma}$ that contains $v$ for $\sigma\sim\mu_{\mathcal{M}}$ some random assignment that we locally construct using depth first search starting from $v$. We initialize $\sigma=\emptyset$ and update $\sigma$ along the way by sampling any marked variables we encounter. Thus, we are able to locally construct the relevant portion of a desired, consistent partial assignment. Shattering results (Lemma 2.11) will give us a polylogarithmic-sized component in $\mathsf{Conn}$ and thus polylogarithmic running time of both $\mathsf{Conn}$ and the subsequent call to $\mathsf{UniformSample}$ (Theorem 4.3). We make this analysis more precise in subsequent sections.
4 Analysis
We verify the correctness and efficiency of Algorithm 1 by analyzing each of its steps. The $\mathsf{IsMarked}$ subroutine has been discussed in Section 2 (Theorem 2.12). In this section, we finish the analysis of $\mathsf{MarginSample}$, $\mathsf{Conn}$ and $\mathsf{UniformSample}$.
4.1 Correctness
Conditioned on $\mathsf{IsMarked}$ returning a consistent marking $\mathcal{M}$ over all $v\in V$ (which by Theorem 2.12 occurs with probability $1-1/n$), we show that with probability 1, $(\mathsf{MarginSample}(v))_{v\in\mathcal{M}}$ follows the marginal distribution $\mu_{\mathcal{M}}$ and computes a variable sample in expected $\mathrm{polylog}(n)$ time.
Correctness of the $\mathsf{MarginSample}$ and $\mathsf{MarginOverflow}$ routines will extend by an identical argument to the corresponding algorithms in [16], since we condition on a consistent marking and compute the same connected component but locally. More precisely, we apply the following results.
Theorem 4.1 ([16, Theorem 5.5]).
We have the following pair of correctness results:
  1. (1)
    Suppose that $(\Phi,\sigma,v)$ satisfies the following trio of conditions:
    • $\Phi=(V,\mathcal{C})$ is a $(k,d)$-formula where $v\in V$ and $\sigma$ is a feasible partial assignment;
    • For every clause $C\in\mathcal{C}$,
      $\mathbb{P}_{\mu}[\text{$C$ violated}\mid\sigma]\leq 2\xi,$
      for $\xi$ defined in 3;
    • The variable $v$ is not part of partial assignment $\sigma$ (i.e., $v\notin\Lambda(\sigma)$).
    Then, $\mathsf{MarginSample}(\Phi,\sigma,v)$ terminates with probability $1$, and it returns a random value $x\in\{0,1\}$ distributed as $\mu_{v}^{\sigma}$ upon termination.
  2. (2)
    Suppose that $(\Phi,\sigma,v)$ satisfies the following trio of conditions:
    • $\Phi=(V,\mathcal{C})$ is a $(k,d)$-formula where $v\in V$ and $\sigma$ is a feasible partial assignment;
    • For every clause $C\in\mathcal{C}$,
      $\mathbb{P}_{\mu}[\text{$C$ violated}\mid\sigma]\leq 2\xi,$
      for $\xi$ defined in 3;
    • The variable $v$ is not part of partial assignment $\sigma$ and furthermore has been given assignment $v\leftarrow\star$.
    Then, $\mathsf{MarginOverflow}(\Phi,\sigma,v)$ terminates with probability $1$, and it returns a random value $x\in\{0,1\}$ distributed as the overflow distribution $\mathcal{D}:=\frac{\mu_{v}^{\sigma}-\theta}{1-2\theta}$ upon termination.
With Theorem 4.1, we can verify the correctness and consistency of $\mathsf{MarginSample}(\Phi,\emptyset,v)$ for each $v\in\mathcal{M}$.
Proposition 4.2.
Let $\Phi$ be a $(k,d)$-formula and $\alpha\in(0,1)$ be a constant such that $d,\alpha$ satisfy the conditions of Theorem 2.12, and
$2^{4}\cdot d^{5}\cdot k^{6}\cdot 2^{-\alpha k}\leq\frac{1}{150e^{3}}.$ (4)
Given a public string of random bits $\mathbf{R}$, suppose that $(\mathsf{IsMarked}(\Phi,v))_{v\in V}$ is an oblivious LCA and gives a valid $\alpha$-marking $\mathcal{M}\subseteq V$. Then, $(\mathsf{MarginSample}(\Phi,\emptyset,v))_{v\in\mathcal{M}}$ is memory-less and follows the distribution $\mu_{\mathcal{M}}$.
@@ 4.1 Correctness
Consider some public randomness $\mathbf{R}$. This determines a global ordering of variables and induces a marking given by $(\mathsf{IsMarked}(v))_{v\in V}$. We condition on the outcome that $(\mathsf{IsMarked}(v))_{v\in V}$ gives a valid $\alpha$-marking $\mathcal{M}\subseteq V$, where =$\mathcal{M}=\{v_{1},\ldots,v_{s}\}$ with $v_{1}<\cdots<v_{s}$ with respect to the random global ordering.
Consider any marked variable $v\in\mathcal{M}$. To sample the assignment of $v$, $\mathsf{MarginSample}(\Phi,\emptyset,v)$ and subsequent recursive calls to $\mathsf{MarginOverflow}$ only attempt to sample the assignments of marked variables encountered along the algorithm. Thus, for any clause $C\in\mathcal{C}$, we have at least $\alpha k$ unmarked variables and thus
$\mathbb{P}_{\mu}[\text{$C$ violated}\mid\sigma]\leq 2^{-\alpha k}\leq 2\xi,$
which holds by our choice of $\xi$ in 3 and $d,\alpha$ in 4. This guarantees that in every call to $\mathsf{MarginSample}$ and $\mathsf{MarginOverflow}$, the conditions for both properties in Theorem 4.1 are satisfied.
We now show consistency of the algorithm, i.e, the outputs $(\mathsf{MarginSample}(\Phi,\emptyset,v))_{v\in\mathcal{M}}$ jointly follow the distribution $\mu_{\mathcal{M}}$. Whenever we compute the assignment of some $v\in\mathcal{M}$, we recompute the assignments of all variables $w$ whose assignment $v$ depends on. Computing the assignment of $w$ involves two steps: (a) first checking where $r_{w}$ falls relative to $\theta_{w}$, (b) calling $\mathsf{MarginOverflow}$ if we cannot immediately assign $w$. Step (a) only relies on the bits of randomness corresponding to $r_{w}$ from the shared randomness $\mathbf{R}$ and is independent of the order in which the variables are processed. Regarding (b), the global ordering and consistency of the marking (Theorem 2.12) ensure that variables are processed in a consistent order and thus retains the desired obliviousness property of the local sampler. ∎
Given a feasible partial assignment $\sigma$ to a satisfiable $(k,d)$-formula $\Phi$, $\mu_{v}^{\sigma}$ is independent of any variable that does not appear in $\Phi_{v}^{\sigma}$ (defined in 2.3), the connected component of the dependency graph containing $v$ after reducing $\Phi$ under partial assignment $\sigma$. The existence of a fast sampling algorithm, $\mathsf{UniformSample}(\cdot)$, for a uniformly random satisfying assignment to $\Phi_{v}^{\sigma}$ follows from the following theorem, which we specialize to our setting.
Theorem 4.3 ([16, Theorems 5.1 and 6.1]).
Suppose $\Phi$ is a $(k,d)$-CNF formula on $n$ variables with $k\cdot 2^{-k}\cdot(dk)^{5}\cdot 4\leq\frac{1}{150e^{3}}$. Then there exists an algorithm $\mathsf{UniformSample}$ that terminates in $O(k^{3}(dk)^{9}n)$ time in expectation, and outputs a uniformly random satisfying assignment for $\Phi$.
Correctness for Algorithm 1 follows from the above correctness of $\mathsf{MarginSample}$ (Proposition 4.2) and $\mathsf{UniformSample}$ (Theorem 4.3) algorithms, along with the independence of marginal distributions in different connected components computed by $\mathsf{Conn}$.
4.2 Efficiency
In this section, we discuss the efficiency of Algorithm 1. We start with the following proposition that gives an upper bound on the size of $\mathsf{Conn}(\Phi,\sigma,v)$ in Algorithm 3, where $v$ is a marked variable.
Proposition 4.4.
Suppose Algorithm 3 is called during the execution of $\mathsf{MarginSample}(\Phi,\emptyset,v)$ with $\mathsf{IsMarked}(v)=1$. Let $\Phi^{\sigma}_{v}=\mathsf{Conn}(\Phi,\sigma,v)$ in line 9 of Algorithm 3. Suppose $d\leq 2^{\alpha k/4}$. Then with probability at least $1-n^{-0.1\log n}$ the number of clauses in $\Phi^{\sigma}_{v}$ is at most $kd\log^{2}n$.
Proof.
Let $\Psi=(V_{\Psi},\mathcal{C}_{\Psi})$ denote the induced $k$-CNF obtained by taking all the clauses in $\Phi$ in their entirety that appear (possibly in a reduced form) in $\Phi^{\sigma}_{v}$. In particular $\Psi$ is $k$-uniform (although $\Phi^{\sigma}_{v}$ might not be), and $\Phi_{v}^{\sigma}$ and $\Psi$ have the same number of clauses. Let $\mathcal{M}_{\Psi}$ be the set of marked variables in $V_{\Psi}$.
For any $T\subseteq\mathcal{M}_{\Psi}$, we say that $T$ is a 2-tree if both of the following conditions are satisfied:
@@ 4.2 Efficiency
    for all $u,v\in T$, $\text{dist}_{H_{\Phi}}(u,v)\geq 2$ (where $H_{\Phi}$ is the dependency hypergraph of 2.1);
@@
    if one adds an edge between every $u,v$ with $\text{dist}_{H_{\Phi}}(u,v)=2$, then $T$ is connected.
@@
Suppose $\Psi$ has $b$ clauses. Then by a greedy removal argument, one can find a 2-tree $T\subseteq\mathcal{M}_{\Psi}$ with $|T|=\left\lfloor\frac{b}{kd}\right\rfloor$. For every $w\in T$, let $C_{w}\in C_{\Psi}$ be any clause that contains $w$. We know that clauses in $\{C_{w}:w\in T\}$ have pairwise disjoint variable sets as $\text{dist}_{H_{\Phi}}(u,v)\geq 2$ for all $u,v\in T$.
@@
Since $\Phi_{v}^{\sigma}$ is a single connected component, we know that for every marked variable $w^{\prime}\in\text{vbl}(C_{w})$, $\sigma(w^{\prime})$ does not satisfy $C_{w}$. This occurs with probability $1-\theta$, as we might set $w^{\prime}\leftarrow\star$ which occurs with probability $1-2\theta$, or we might set $w^{\prime}\leftarrow x\in\{0,1\}$ such that $(C_{w})_{w^{\prime}\leftarrow x}\neq 1$, which occurs with probability $\theta$.
Recall that every $C_{w}$ has at least $\alpha k$ marked variables. Moreover, all of these variables have been accessed by the partial assignment $\sigma$, as we call $\mathsf{MarginOverflow}$ on marked variables in a depth first search fashion. Thus, we get that
$\mathbb{P}[\text{$C_{w}$ is not satisfied by $(\sigma(w^{\prime}))_{w^{\prime}\in\text{vbl}(C_{w})\cap\mathcal{M}}$}]\leq(1-\theta)^{\alpha k}.$
@@
Since the $C_{w}$ have disjoint variable sets, the events whether $C_{w}$ are satisfied by $(\sigma(w^{\prime}))_{w^{\prime}\in\text{vbl}(C_{w})\cap\mathcal{M}}$ are independent of each other. Thus, we know that
@@
$\mathbb{P}[\text{none of the $C_{w}$, $w\in T$ is satisfied by $\sigma$}]\leq(1-\theta)^{\alpha k|T|}=(1-\theta)^{\alpha k\left\lfloor\frac{b}{kd}\right\rfloor}.$
@@
Finally, we union bound over all 2-trees of size $\left\lfloor\frac{b}{kd}\right\rfloor$ that contains $v$. It is well-known (see [8, Corollary 5.7]) that, since $H_{\Phi}$ is a $k$-uniform hypergraph with maximum degree $d$, the number of 2-trees $T\subseteq H_{\Phi}$ containing $v$ with $|T|=\ell$ is at most $\frac{(ed^{2})^{\ell-1}}{2}$. Setting $\ell=\left\lfloor\frac{b}{kd}\right\rfloor$ and applying a union bound, we get that
@@
$\mathbb{P}[\text{$\Phi_{v}^{\sigma}$ has $b$ clauses}]=\mathbb{P}[\text{$\Psi$has $b$ clauses}]\leq\frac{(ed^{2})^{\left\lfloor\frac{b}{kd}\right\rfloor-1}}{2}(1-\theta)^{\alpha k\left\lfloor\frac{b}{kd}\right\rfloor}\leq 0.9^{\left\lfloor\frac{b}{kd}\right\rfloor}.$

### v2

@@ Appendix B Efficiency of Algorithm 6
Let $\Psi=(V_{\Psi},\mathcal{C}_{\Psi})$ be a $k$-CNF, $\sigma$ be a feasible partial assignment of $\Psi$, and $\Psi^{\sigma}=(V_{\Psi}^{\sigma},\mathcal{C}_{\Psi}^{\sigma})$ be the reduced CNF (see 2.4). Suppose we have
[$\mathbb{P}[\lnot C\mid\sigma]\leq\zeta\qquad\text{for every $C\in\mathcal{C}_{\Psi}^{\sigma}$}.$
Then there exists an algorithm $\mathsf{BF}(\cdot)$ such that for every $v\in V^{\sigma}$,
  • $\mathsf{BF}(v)$ with probability 1 returns $\xi\sim(\mu_{\Psi})_{v}^{\mathrm{pad},\sigma}$;
  • $\mathsf{BF}(v)$ has expected cost $O(k^{9}d^{6}|\mathcal{C}_{\Psi}^{\sigma}|(1-e\zeta)^{-|\mathcal{C}_{\Psi}^{\sigma}|})$.
We remark that in Algorithm 6, our partial assignment $\sigma_{\Lambda}$ is always supported on the marked variables. Since every clause has at least $\alpha k$ unmarked variables that are not assigned by $\sigma_{\Lambda}$, we get that [$\mathbb{P}[\lnot C\mid\sigma_{\Lambda}]\leq 2^{-\alpha k}$ for every $C$. Thus when applying Proposition B.2, we will take $\zeta=2^{-\alpha k}$.
We can now give an upper bound on the expected cost of Algorithm 6, proving Lemma 4.7.
Proof of Lemma 4.7.
[…]
Observe that for every $t^{\prime}$, if $\mathsf{Glauber}(t^{\prime})$ has been computed once, then we will permanently assign $M(t^{\prime})=\mathsf{Glauber}(t^{\prime})\in\{0,1\}$, and later executions of $\mathsf{Glauber}(t^{\prime})$ will terminate at Line 4. Thus, it suffices to upper bound the cost of every first execution of $\mathsf{Glauber}(t^{\prime})$ before entering another $\mathsf{Glauber}(t^{\prime\prime})$. Multiplying this by $|A|$ would give an upper bound on the cost of $\mathsf{Glauber}(t)$.
Suppose we are at the first time of executing $\mathsf{Glauber}(t^{\prime})$ for some $t^{\prime}$. We first estimate the cost of the while loop in Lines 10–18, which is a constant multiple of the number of executions of Line 14 and Line 18. Note that every time we execute Line 14 or Line 18, some $w$ is added to $\Lambda$ due to a clause $C$ chosen in Line 11, with $w\in\operatorname{vbl}(C)$. Moreover, each clause $C$ can be chosen in Line 11 at most once. Thus, the number of pairs $(w,C)$ that correspond to an execution of Line 14 or Line 18 is at most $d|\Lambda|$. Consequently, the cost of the while loop in Lines 10–18 is $O(d|\Lambda|)=O(kd^{3}|R|/\alpha)$.
We now estimate the cost of Line 19. Observe that in Line 19, we can apply the Bernoulli factory in Proposition B.2 to compute the padding distribution $(\nu_{\Psi})_{u}^{\mathrm{pad},\sigma_{\Lambda}}$. Since the connected component $\Psi$ in Line 19 has at most $d|V^{\prime}|\leq 2d^{2}|R|/\alpha$ clauses, by Proposition B.2, the expected cost of Line 19 is at most
$O(k^{9}d^{8}|R|/\alpha(1-e2^{-\alpha k})^{-2d^{2}|R|/\alpha}).$
Let $R_{\max}=80dk^{4}\log n$. Combining the above and applying 5, we get that the expected cost of $\mathsf{Glauber}(t)$ is at most
$\mathbb{E}[|A|\cdot O(kd^{3}|R|/\alpha+k^{9}d^{8}|R|/\alpha(1-e2^{-\alpha k})^{-2d^{2}|R|/\alpha})]$
$\leq\mathbb{E}[|R|\cdot O(kd^{3}|R|/\alpha+k^{9}d^{8}|R|/\alpha(1-e2^{-\alpha k})^{-2d^{2}|R|/\alpha})]$
$\leq\sum_{r=0}^{R_{\max}}\mathbb{P}[|R|\geq r]\cdot r\cdot O(kd^{3}r/\alpha+k^{9}d^{8}r/\alpha(1-e2^{-\alpha k})^{-2d^{2}r/\alpha})$
$\leq\sum_{r=0}^{R_{\max}}O(2^{-\frac{r}{48dk^{4}}}\cdot r\cdot kd^{3}r/\alpha+k^{9}d^{8}r/\alpha(1-e2^{-\alpha k})^{-2d^{2}r/\alpha})$
$\leq\sum_{r=0}^{R_{\max}}O\left(2^{-\frac{r}{48dk^{4}}}\cdot r\cdot\left(\frac{kd^{3}r}{\alpha}+\frac{k^{9}d^{8}r}{\alpha}\cdot e^{\frac{2d^{2}r/\alpha}{2^{\alpha k}}}\right)\right)=O(k^{9}d^{8}R_{\max}^{2}/\alpha)=O(k^{17}d^{10}\log^{2}n/\alpha),$
where in the last step we used the information that
$2^{-\frac{1}{48dk^{4}}}\cdot e^{\frac{2d^{2}/\alpha}{2^{(1-\alpha)k}}}\leq 0.9\quad\Rightarrow\quad\sum_{r=0}^{R_{\max}}\left(2^{-\frac{1}{48dk^{4}}}\cdot e^{\frac{2d^{2}/\alpha}{2^{\alpha k}}}\right)^{r}=O(1).$
Appendix C Proof of Theorem 3.3
In this section we show Theorem 3.3. Let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking. We introduce the local access algorithm $\mathsf{Conn}$ (Algorithm 7) such that for every unmarked variable $v\in V\setminus\mathcal{M}$, $\mathsf{Conn}(v)$ returns the connected component of $\Phi^{(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}}$ that contains $v$.
Algorithm 7 $\mathsf{Conn}(v)$
Input: CNF formula $\Phi=(V,\mathcal{C})$, $v\in V\setminus\mathcal{M}$
      Output: random assignment in $\{0,1\}$
1:Initialize $\mathcal{S}=\{C\in\mathcal{C}\mid v\in\text{vbl}(C)\}$ as a stack
@@
6:              $\sigma(w)\leftarrow\mathsf{MarginSample}(w)$
@@
12:              Push $\{C^{\prime}\in\mathcal{C}\mid w\in\operatorname{vbl}((C^{\prime})^{\sigma}),(C^{\prime})^{\sigma}\neq 1,C^{\prime}\text{ has never been added to $\mathcal{S}$}\}$ onto $\mathcal{S}$               return $\Phi_{v}^{\sigma}$
For every clause $C\in\mathcal{C}$ and partial assignment $\sigma$, we define $C^{\sigma}$ to be the reduced clause after conditioning on $\sigma$ and removing the variables assigned 0 or 1 by $\sigma$ (in particular, $C^{\sigma}=1$ if $C$ is already satisfied by $\sigma$). It is not hard to read from the algorithm that $\mathsf{Conn}(v)$ computes the connected component of $\Phi^{(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}}$ that contains $v$. Starting with the empty assignment $\sigma=\emptyset$ and $\mathcal{S}$ to be the clauses that contain $v$, the algorithm keeps sampling the marked variables touched by $\mathcal{S}$, adding these assignments to $\sigma$, and see if clauses in $\mathcal{S}$ are satisfied by $\sigma$. If a clause is satisfied by $\sigma$, then we simply remove it from $\mathcal{S}$; if it is not satisfied by $\sigma$, then in addition to removing it from $\mathcal{S}$, we also add all unexplored clauses adjacent to this clause to $\mathcal{S}$. One can see that in this procedure, $\mathcal{S}$ maintains to be the boundary of those explored and yet unsatisfied clauses given by the partial assignment $\sigma$. Once $\mathcal{S}$ becomes empty, it means that in the reduced CNF $\Phi^{\sigma}$, there is no unexplored clause adjacent to the explored clauses that are not satisfied by $\sigma$. In other words, $\sigma$ has shattered $\Phi$ so that those explored clauses form a connected component $\Phi_{v}^{\sigma}$.
To show Theorem 3.3, it remains to show that the algorithm $\mathsf{Conn}(v)$ has expected cost $\mathrm{poly}\log(n)$. We show this by giving an upper bound on the size of the connected component of $\Phi^{(\mathsf{MarginSample}(v))_{u\in\mathcal{M}}}$ that contains $v$.
Proposition C.1.
Fix $v\in V\setminus\mathcal{M}$. Let $\tau=(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$, and let $\Phi_{v}^{\tau}$ denote the connected component of $\Phi^{\tau}$ that contains $v$. Suppose $\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq 0.4$ and $d\leq 2^{\alpha k/4}$. Then with probability at least $1-n^{-0.1\log n}$ the number of clauses in $\Phi^{\tau}_{v}$ is at most $kd\log^{2}n$.
@@
Let $\Psi=(V_{\Psi},\mathcal{C}_{\Psi})$ denote the induced $k$-CNF obtained by taking all the clauses in $\Phi$ in their entirety that appear (possibly in a reduced form) in $\Phi^{\tau}_{v}$. In particular $\Psi$ is $k$-uniform (although $\Phi^{\tau}_{v}$ might not be), and $\Phi_{v}^{\tau}$ and $\Psi$ have the same number of clauses. Let $\mathcal{M}_{\Psi}$ be the set of marked variables in $V_{\Psi}$.
Recall the dependency hypergraph $H_{\Phi}=(V,\mathcal{E})$ defined in 2.2. Recall also from Definition A.3 that $Z\subseteq V$ is a 2-tree if both of the following conditions are satisfied:
@@
    for all $u,v\in Z$, $\text{dist}_{H_{\Phi}}(u,v)\geq 2$;
@@
    the associated graph with vertex set $Z$ and edge set $\{\{u,v\}\subseteq Z:\mathrm{dist}_{H_{\Phi}}(u,v)=2\}$ is connected.
@@
Suppose $\Psi$ has $b$ clauses. Then by a greedy removal argument, one can find a 2-tree $Z\subseteq\mathcal{M}_{\Psi}$ with $|Z|=\left\lfloor\frac{b}{kd}\right\rfloor$. For every $w\in Z$, let $C_{w}\in C_{\Psi}$ be any clause that contains $w$. We know that clauses in $\{C_{w}:w\in Z\}$ have pairwise disjoint variable sets as $\text{dist}_{H_{\Phi}}(u,v)\geq 2$ for all $u,v\in Z$.
@@
Since $\Phi_{v}^{\tau}$ is a single connected component, we know that for every marked variable $w^{\prime}\in\text{vbl}(C_{w})$, $\tau(w^{\prime})$ does not satisfy $C_{w}$. This occurs with probability at most $1-\theta$, as we know that $\nu_{\mathcal{M}}$ is $\theta$-lower bounded, and $\mathsf{MarginSample}(w^{\prime})$ returns each of 0 and 1 with probability at least $\theta$.
Since every $C_{w}$ has at least $\alpha k$ marked variables, we get that
$\mathbb{P}[\text{$C_{w}$ is not satisfied by $(\tau(w^{\prime}))_{w^{\prime}\in\text{vbl}(C_{w})\cap\mathcal{M}}$}]\leq(1-\theta)^{\alpha k}.$
@@
Since the $C_{w}$ have disjoint variable sets, the events whether $C_{w}$ are satisfied by $(\tau(w^{\prime}))_{w^{\prime}\in\text{vbl}(C_{w})\cap\mathcal{M}}$ are independent of each other. Thus, we know that
@@
$\mathbb{P}[\text{none of the $C_{w}$, $w\in Z$ is satisfied by $\tau$}]\leq(1-\theta)^{\alpha k|Z|}=(1-\theta)^{\alpha k\left\lfloor\frac{b}{kd}\right\rfloor}.$
@@
Finally, we union bound over all 2-trees of size $\left\lfloor\frac{b}{kd}\right\rfloor$ that contains $v$. Recall from A.4 that, since $H_{\Phi}$ is a hypergraph with maximum degree $d$, the number of 2-trees $Z\subseteq H_{\Phi}$ containing $v$ with $|Z|=\ell$ is at most $\frac{(ed^{2})^{\ell-1}}{2}$. Setting $\ell=\left\lfloor\frac{b}{kd}\right\rfloor$ and applying a union bound, we get that
@@
$\mathbb{P}[\text{$\Phi_{v}^{\tau}$ has $b$ clauses}]=\mathbb{P}[\text{$\Psi$ has $b$ clauses}]\leq\frac{(ed^{2})^{\left\lfloor\frac{b}{kd}\right\rfloor-1}}{2}(1-\theta)^{\alpha k\left\lfloor\frac{b}{kd}\right\rfloor}\leq 0.9^{\left\lfloor\frac{b}{kd}\right\rfloor}.$